OpenStudy (anonymous):

Divide the complex number: 7i/8+i (i= imaginary #) Show work?

OpenStudy (anonymous):

Multiply top and bottom by the conjugate: \(\frac{7i}{8+i}\times\frac{8-i}{8-i}=\frac{56i+7}{65}=7+\frac{56}{65}i\).

OpenStudy (anonymous):

Thanks ^_^

OpenStudy (anonymous):

You're welcome! :)

jimthompson5910 (jim_thompson5910):

\[\Large \frac{7i}{8+i}\] \[\Large \frac{(7i)(8-i)}{(8+i)(8-i)}\] \[\Large \frac{(7i)(8-i)}{(8+i)(8-i)}\] \[\Large \frac{(7i)(8)+(7i)(-i)}{(8+i)(8-i)}\] \[\Large \frac{(7i)(8)+(7i)(-i)}{(8)(8)+(8)(-i)+(i)(8)+(i)(-i)}\] \[\Large \frac{56i-7i^2}{64-8i+8i-i^2}\] \[\Large \frac{56i-7i^2}{64-8i+8i-(-1)}\] \[\Large \frac{56i-7i^2}{64-8i+8i+1}\] \[\Large \frac{7+56i}{65}\] \[\Large \frac{7}{65}+\frac{56}{65}i\] \[\Large So\frac{7i}{8+i}=\frac{7}{65}+\frac{56}{65}i\]

OpenStudy (anonymous):

Oh sorry I made a typo in the last step. Nice work and Latex Jim!!